Opened 7 years ago

Closed 3 years ago

#445 closed flight processing (fixed)

EU12/12, flight day 141b/2012, Santorini

Reported by: knpa Owned by: adbe
Priority: alpha 4 high Milestone: 2012 data processing completion
Component: Processing: general Keywords:
Cc: Other processors:

Description (last modified by adbe)

Data location: /users/rsg/arsf/arsf_data/2012/flight_data/greece/EU12_12-2012_141b_Santorini

Data arrived from Greece via SATA disk on 13/07/2012

EU12/12 Scientific objective:

EU12/12 Priority: a4m

PI: David Pyle

Sensors:

Camera (17/08/2012)
Eagle (30/08/2012)
Hawk (30/08/2012)
LiDAR (17/08/2012)

Change History (17)

comment:1 Changed 7 years ago by knpa

One of the lidar flightlines (at 1326) is broken up into separate raw directories.

Last edited 7 years ago by knpa (previous) (diff)

comment:2 Changed 7 years ago by adbe

Processed Navigation.

comment:3 Changed 7 years ago by adbe

Delivery created for Camera and LiDAR

comment:4 Changed 7 years ago by anhi

Last edited 7 years ago by anhi (previous) (diff)

comment:5 Changed 7 years ago by knpa

Starting Camera delivery check.

comment:6 Changed 7 years ago by knpa

No problems found, RCD ready to deliver.

comment:7 Changed 7 years ago by emca

Started to delivery check the lidar before I realised it was already done.
Two things I noticed:

  • The handwritten logsheet in the delivery should be removed
  • Logsheet should mention in the notes column where there is no Eagle/Hawk data for those lines (it's not relevant for the lidar delivery but just so that we have a consistent logsheet for both deliveries)

comment:8 Changed 7 years ago by emca

Lidar delivery check is not done (comment 4 above was entered in wrong ticket).
Restarting delivery check.

comment:9 Changed 7 years ago by emca

Lidar delivery check complete.

Readme - height differences contains line names that don't exist (124956 + 132646)
Also the above comments re: logsheet

No issues with the data. Otherwise ready to delivery.

comment:10 Changed 7 years ago by adbe

  • Description modified (diff)

Camera and LiDAR delivered to David Pyle on 17/08/2012 on disk 121.

comment:11 Changed 7 years ago by adbe

Hyperspectral delivery created.

comment:12 Changed 7 years ago by anhi

Hyperspectral delivery check complete

  • Eagle line 10 level three has dropped scan, mention in Readme
  • Eagle line 4 not aligned in the mosaic
  • Some dropped scans in level ones: line e10 and e11
  • Zipped map files need creating

comment:13 Changed 7 years ago by adbe

Made changes to hyperspectral delivery

comment:14 Changed 7 years ago by adbe

  • Description modified (diff)

Delivered Hyperspectral to David Pyle on disk 128 on 30/08/2012

comment:15 Changed 6 years ago by stgo

Beginning archiving.

comment:16 Changed 4 years ago by dac

Archiving

Uploading data to NEODC

comment:17 Changed 3 years ago by dac

  • Resolution set to fixed
  • Status changed from new to closed

Archiving

Raw and processed data now available from NEODC.

Note: See TracTickets for help on using tickets.